Question 16 (1 point)

The local government in Central City is considering using an alternative to a Tax Deferred Annuity as a retirement plan. Which of the following could Central City Government use?

Question 16 options:

401(k) plan

SIMPLE IRA

SIMPLE 401(k)

Section 403(b)

city governments can only use a Tax Deferred Annuity

Question 17 (1 point)

Donor made a gift of property to donee. At the time of the gift, the donor had a $90,000 basis in the property, the property had a fair market value equal to $80,000, and the donor paid gift tax of $10,000 with respect to the gift. If the donee sells the property for $75,000, the donee's basis for purposes of determining loss is equal to

Question 17 options:

$75,000

$80,000

$90,000

$100,000

Question 18 (1 point)

With regard to post mortem income tax elections all of the following are true if the decedent transferred property before death, except:

Question 18 options:

If a decedent made an installment sale, the estate can elect of of installment reporting

If property of the decedent was involuntarily converted (i.e., destroyed by natural disaster, condemned) the estate or successor may be able to replace it and avoid taxation of gain to decedent for income tax purposes

A decedent must be declared of sound mind before transferring any property before death

A decedent has the right to make transfers of property before death

Question 19 (1 point)

Maria Valquez is a public school teacher. Her employer provides a Tax Deferred Annuity (TDA). She began working for this employer four years ago and started her TDA at that time. Over those four years, she has contributed $1000, $2500, $3000, $3000 to her TDA through salary reduction. Her employer matches $1 for $1 up to $100 and offers graded vesting at the rate required by law for TDA accounts. Currently, Maria’s vested interest in the plan is:

Question 19 options:

$9,900

$9,740

$9,660

$9,500

$9,580

Question 20 (1 point)

Sam, a U.S. citizen, died in 2011. At his death, he owned real estate in Canada and in Ohio. Sam received the Ohio real estate 15 years earlier from his brother's estate. The brother's estate had paid federal estate tax on the Ohio real estate. Assume Sam's estate paid death taxes to Canada and Ohio. A deduction or credit is not available to Sam for federal estate tax purposes for which of the following?

Question 20 options:

State death taxes paid to Ohio

Foreign death taxes paid to Canada

Any of the estate tax paid by Sam's brother on the prior transfer.

The unified credit
Answer #1

Dear Student,

As per the Chegg Policy, only the first four questions should be answered. Kindly take note of it.

Part 16

Option B Simple IRA

As it is simple to understand and helps small employers to contribute toward their employees' as well as their own retirement savings, it is recommended to be used as an alternative to a Tax Deferred Annuity as a retirement plan.

Part 17

Option B $80,000

At the time of the gift, if the property is a loss property that is donor’s basis > fair market value, then the donee has to use the lower value of donor’s basis or fair market value for the purpose of calculating loss associated with the disposition of asset. This condition applies only for calculating loss and not for the purpose of calculating gain.

Part 18

Option B

If property of the decedent was involuntarily converted (i.e., destroyed by natural disaster, condemned) the estate or successor may be able to replace it and avoid taxation of gain to decedent for income tax purposes

Part 19

Option B 9740

Tax Deferred Annuity contribution = 1000+2500+3000+3000 = $9500

Maria’s employer contribution = 400

Maria’s vested amount = 400*60% = 240 (as she worked for 4 years)

Therefore,

Maria’s total vested interest = 9500+240 = 9740